# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ 2.6 Update
After performing the previous phases, the live network $\mathcal{L}$ must be coupled with the newly updated clean network $\mathcal{C}$. A high level description on how to perform such update in constant time is the following: during the duplication phase a “snapshot” (e.g., a copy) of the clean network $\mathcal{C}$ is taken and used as the new $\mathcal{L}$. Intuitively, this procedure requires $\mathcal{O}(1)$ rounds because each node in $\mathcal{C}$ is taking care of the snapshot that can be considered as local computation. However, we must be careful with the notion of snapshot. To preserve dynamic resource-competitiveness we show how to avoid creating new edges while taking a snapshot of $\mathcal{C}$ during this phase. Assume that every edge in the clean network $\mathcal{C}$ has a length two local [^11] Meaning that each node $u\in\mathcal{C}$ has its own copy of the label $\lambda$ for each edge incident to it. Observe that for an edge $(u,v)$ the local label of $u$ can not differ from the one of $v$. binary label $\lambda$ in which the first coordinate indicates its presence in $\mathcal{C}$ and the second one in $\mathcal{L}$. Moreover, each label $\lambda$ can be of three different types: (1) “$10$” the edge is in $\mathcal{C}$ not in $\mathcal{L}$; (2) “$01$” the edge is not in $\mathcal{C}$ but it is in $\mathcal{L}$; and, (3) “$11$” the edge is in both $\mathcal{C}$ and $\mathcal{L}$. Moreover, assume that each node $u$ in $\mathcal{C}$ has the label $\lambda(u,v)$ associated to its port encoding the connection with node $v$. Then, each node $u$ in $\mathcal{C}$ for each $0\leq\ell\leq\ell_{\max}(u)$ performs the following local computation:
@@
If $u$’s port-label associated to the edge connected to the left neighbor at level $\ell$ (i.e., $\lambda(u,v)$ such that $v\in N_{\mathcal{C}}^{(L,\ell)}(u)$) is “$10$” then set it to “$11$”;
@@
If $u$’s port-label associated to the edge connected to the right neighbor at level $\ell$ (i.e., $\lambda(u,w)$ such that $w\in N_{\mathcal{C}}^{(R,\ell)}(u)$) is “$10$” then set it to “$11$”;
Notice that during this phase no edge has the label “$01$” because $\mathcal{C}$ was “cleaned” during the deletion phase. All the nodes in $\mathcal{C}$ executed the labeling procedure, and the live network $\mathcal{L}$ is given by the edges with label “$11$”. It follows that this phase requires a constant number of rounds.
@@
Proof 2.23.
In order to prove that our maintenance protocol is dynamic resource competitive we suffice to show that a generic iteration of our maintenance cycle respects this invariant.
@@
We conclude by showing that the skip list is maintained with high probability for $n^{c}$ rounds, where $c\geq 1$ is an arbitrary big constant.
@@
Proof 2.25.
Each phase of the maintenance protocol (i.e., Algorithm 1) succeeds with probability at least $1-\frac{1}{n^{c}}$, for some arbitrary big constant $c\geq 1$. Let $X$ be a geometric random variable of parameter $p=\frac{1}{n^{c}}$ that counts the number of cycle needed to have the first failure, then its expected value is $\mathbb{E}\left[X\right]=n^{c}$, for $c\geq 1$. Moreover, the probability of maintenance protocol succeeding in a round $n^{r}$ for some $r<c$ is $(1-\frac{1}{n^{c}})^{n^{r}}$, Without loss of generality assume $c>1$ and $r=c-1$, then $(1-\frac{1}{n^{c}})^{n^{c-1}}\leq e^{-n^{c-1}/n^{c}}\geq 1-\frac{1}{n}$.
@@
To prove the main theorem it is sufficient to notice that the skip list will always be connected and will never loose its structure thanks to Lemma 2.1. Indeed, every time a group of nodes is removed from the network, there are committees that in $\mathcal{O}(1)$ rounds will take over and act on their behalf. This allows all the queries to go through without any slowdown. Thus all the queries will be executed in $\mathcal{O}(\log n)$ rounds with high probability even with a churn rate of $\mathcal{O}(n/\log n)$ per round. Next we can show that the distributed data structure can be efficiently maintained using maintenance cycles of $\mathcal{O}(\log n)$ rounds each. To do this, it is sufficient to show that each phase of our maintenance cycle (Algorithm 1) can be carried out in $\mathcal{O}(\log n)$ rounds (see Lemma 2.3, Lemma 2.8, Lemma 2.19, and Lemma 2.21). Moreover, from Lemma 2.22 we have that each phase is $(\alpha,\beta)$-dynamic resource competitive with $\alpha=\mathcal{O}(\log n)$ and $\beta=\mathcal{O}(\text{polylog}{(n)})$. Finally, the maintenance protocol ensures that the distributed data structure is maintained for at least $n^{c}$ rounds with high probability where $c\geq 1$ is an arbitrarily large constant.
@@ 2.7 Extending our approach to other data structures
The outlined maintenance protocol provides a convenient framework for constructing churn-resilient data structures. Furthermore, it can be easily adapted to maintain a more complex distributed pointer based data structure such as a skip graphs [44, 3, 39, 47]. Indeed, with minor adjustments to the phases outlined in Algorithm 1, a maintenance cycle capable of maintain such data structures against heavy churn rate can be devised. To this end, we briefly discuss how to adapt our results to skip graphs. A skip graph [3], can be viewed as an extension of skip lists. Indeed, both consists of a set of increasingly sparse doubly-linked lists ordered by levels starting at level 00, where membership of a particular node $u$ in a list at level $\ell$ is determined by the first $\ell$ bits of an infinite sequence of random bits associated with $u$, referred to as the membership vector of $u$, and denoted by $m(u)$. Let the first $\ell$ bits of $m(u)$ as $m(u)|\ell$. In the case of skip lists, level $\ell$ has only one list, for each $\ell$, which contains all elements $u$ such that $m(u)|\ell=1^{\ell}$, i.e., all elements whose first $\ell$ coin flips all came up heads. Skip graphs, instead, have $2^{\ell}$ lists at level $\ell$, which we can index from 00 to $12^{\ell}-$ 1. Node $u$ belongs to the $j$-th list of level $\ell$ if and only if $m(u)|\ell$ corresponds to the binary representation of $j$. Hence, each node is present in one list of every level until it eventually becomes the only member of a singleton list. Without loss of generality, assume that the skip graph has sentry nodes as classical skip lists [^12] Thus, at each level $\ell$ there are $2^{\ell}$ sentinels (one for each list in such level).. Thus, the skip graph can be maintained as follows:
@@
To build the base level of the buffer skip graph we can use the same $\mathcal{O}(\log n)$ rounds approach described in Section 2.4. While, to construct the skip graph from level 00, we build an “augmented” skip list in which each node $u$ in each level $\ell$ is identified by its ID/key and the list $j$ at level $\ell$ in which $u$ appears (there are $2^{\ell}$ lists at level $\ell$). Thus, each node $u$ at some level $\ell$ can be: (i) effective in the list $j$ and fill-in for the remaining $i\neq j$, for $j,i\in[1,2^{\ell}]$; or, (ii) fill-in in all the lists $j\in[1,2^{\ell}]$. Finally, such augmented skip list is transformed into a skip graph by running (in parallel) Algorithm 2 in Section 2.3 on each level $\ell$ for which there exists at least one fill-in node.
@@ 2.8 Extending our approach to deal with multiple keys on each node
Throughout the paper we assumed that each node in the network posses exactly one element of the data structure. This assumption can be relaxed to deal with multiple keys on each node. Indeed, in the case in which in the network there are $t\cdot n$ keys, where $t>0$, all the techniques described above allow to maintain the data structure in the presence of the same adversarial churn rate.
@@
Given a network with $n$ nodes in which each vertex $v$ possesses $t$ elements in the skip list for some $t>1$. Then maintenance protocol requires $\mathcal{O}(t\cdot\log n)$ rounds to build and maintain a resilient skip list that can withstand heavy adversarial churn at a churn rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round.
@@ 3 Concluding remarks and discussion
In this work we proposed the first churn resilient skip list that can tolerate a heavy adversarial churn rate of $\mathcal{O}(n/\log n)$ nodes per round. The data structure can be seen as a four networks architecture in which each network plays a specific role in making the skip list resilient to churns and keeping it continuously updated. Moreover, we provided efficient $\mathcal{O}(\log n)$ rounds resource competitive algorithms to (i) delete a batch of elements from a skip list (ii) create a new skip list and, (iii) merge together two skip lists. This last result is the first algorithm that can merge two skip list (as well as a skip list and a batch of new nodes) in $\mathcal{O}(\log n)$ rounds w.h.p.. We point out that these algorithms can be easily adapted to work on skip graphs [3, 39, 47].
In a broader sense, our technique is general and can be used to maintain any kind of distributed data structure despite heavy churn rate. The only requirement is to devise efficient delete, buffer creation, merge, and update algorithms for the designated data structure.
An additional contribution of our work is the improvement on the $\mathcal{O}(\log^{3}n)$ rounds state-of-the-art technique for sorting in the $\text{NCC}_{0}$ model. We show how to sort $n$ elements (despite a high churn rate) in $\mathcal{O}(\log n)$ rounds using results for sorting network theory. However, given the impracticability of the AKS sorting network, our result is purely theoretical. In practice, it could be easily implemented using Batcher’s network instead of the AKS one. This change would slow down the bootstrap phase and the maintenance cycle to $\mathcal{O}(\log^{2}n)$ rounds.

### v2

@@ 2.6 Update
After performing the previous phases, the live network $\mathcal{L}$ must be coupled with the newly updated clean network $\mathcal{C}$. A high level description on how to perform such update in constant time is the following: during the duplication phase a “snapshot” (e.g., a copy) of the clean network $\mathcal{C}$ is taken and used as the new $\mathcal{L}$. Intuitively, this procedure requires $\mathcal{O}(1)$ rounds because each node in $\mathcal{C}$ is taking care of the snapshot that can be considered as local computation. However, we must be careful with the notion of snapshot. To preserve dynamic resource-competitiveness we show how to avoid creating new edges while taking a snapshot of $\mathcal{C}$ during this phase. Assume that every edge in the clean network $\mathcal{C}$ has a length two local [^10] Meaning that each node $u\in\mathcal{C}$ has its own copy of the label $\lambda$ for each edge incident to it. Observe that for an edge $(u,v)$ the local label of $u$ can not differ from the one of $v$. binary label $\lambda$ in which the first coordinate indicates its presence in $\mathcal{C}$ and the second one in $\mathcal{L}$. Moreover, each label $\lambda$ can be of three different types: (1) “1010” the edge is in $\mathcal{C}$ not in $\mathcal{L}$; (2) “0101” the edge is not in $\mathcal{C}$ but it is in $\mathcal{L}$; and, (3) “1111” the edge is in both $\mathcal{C}$ and $\mathcal{L}$. Moreover, assume that each node $u$ in $\mathcal{C}$ has the label $\lambda(u,v)$ associated to its port encoding the connection with node $v$. Then, each node $u$ in $\mathcal{C}$ for each $0\leq\ell\leq\ell_{\max}(u)$ performs the following local computation:
@@
If $u$’s port-label associated to the edge connected to the left neighbor at level $\ell$ (i.e., $\lambda(u,v)$ such that $v\in N_{\mathcal{C}}^{(\texttt{left},\ell)}(u)$) is “1010” then set it to “1111”;
@@
If $u$’s port-label associated to the edge connected to the right neighbor at level $\ell$ (i.e., $\lambda(u,w)$ such that $w\in N_{\mathcal{C}}^{(\texttt{right},\ell)}(u)$) is “1010” then set it to “1111”;
Notice that during this phase no edge has the label “0101” because $\mathcal{C}$ was “cleaned” during the deletion phase. All the nodes in $\mathcal{C}$ executed the labeling procedure, and the live network $\mathcal{L}$ is given by the edges with label “1111”. It follows that this phase requires a constant number of rounds.
@@
Proof 2.25.
In order to prove that our maintenance protocol is dynamic resource competitive it suffices to show that a generic iteration of our maintenance cycle respects this invariant.
@@
We are now ready to show that the skip list is maintained with high probability for $n^{d}$ rounds, where $d\geq 1$ is an arbitrary big constant.
@@
Proof 2.27.
Each phase of the maintenance protocol (i.e., Algorithm 1) succeeds with probability at least $1-\frac{1}{n^{d}}$, for some arbitrary big constant $h\geq 1$. Let $X$ be a geometric random variable of parameter $p=\frac{1}{n^{d}}$ that counts the number of cycle needed to have the first failure, then its expected value is $\mathbb{E}\left[X\right]=n^{d}$, for $d\geq 1$. Moreover, the probability of maintenance protocol succeeding in a round $n^{r}$ for some $r<d$ is $(1-\frac{1}{n^{d}})^{n^{r}}$, Without loss of generality assume $d>1$ and r=d1r=d-1, then $(1-\frac{1}{n^{d}})^{n^{d-1}}\leq e^{-n^{d-1}/n^{d}}\geq 1-\frac{1}{n}$.
We conclude by proving the main theorem in Section 1.2.
@@
To prove the main theorem it is sufficient to notice that the skip list will always be connected and will never lose its structure thanks to Lemma 2.1. Indeed, every time a group of nodes is removed from the network, there are committees that in $\mathcal{O}(1)$ rounds will take over and act on their behalf. This allows all the queries to go through without any slowdown. Thus all the queries will be executed in $\mathcal{O}(\log n)$ rounds with high probability even with a churn rate of $\mathcal{O}(n/\log n)$ per round. Next we can show that the distributed data structure can be efficiently maintained using maintenance cycles of $\mathcal{O}(\log n)$ rounds each. To do this, it is sufficient to show that each phase of our maintenance cycle (Algorithm 1) can be carried out in $\mathcal{O}(\log n)$ rounds (see Lemma 2.3, Lemma 2.8, Lemma 2.21, and Lemma 2.23). Moreover, from Lemma 2.24 we have that each phase is $(\alpha,\beta)$-dynamic resource competitive with $\alpha=\mathcal{O}(\log n)$ and $\beta=\mathcal{O}(\text{polylog}{(n)})$. Finally, the maintenance protocol ensures that the distributed data structure is maintained for at least $n^{c}$ rounds with high probability where $c\geq 1$ is an arbitrarily large constant.
@@ 2.7 Extending our approach to other data structures
The outlined maintenance protocol provides a convenient framework for constructing churn-resilient data structures. Furthermore, it can be easily adapted to maintain a more complex distributed pointer based data structure such as a skip graphs [44, 3, 39, 47]. Indeed, with minor adjustments to the phases outlined in Algorithm 1, a maintenance cycle capable of maintaing such data structures against heavy churn rate can be devised. To this end, we briefly discuss how to adapt our results to skip graphs. A skip graph [3], can be viewed as an extension of skip lists. Indeed, both consists of a set of increasingly sparse doubly-linked lists ordered by levels starting at level 0, where membership of a particular node $u$ in a list at level $\ell$ is determined by the first $\ell$ bits of an infinite sequence of random bits associated with $u$, referred to as the membership vector of $u$, and denoted by $m(u)$. Let the first $\ell$ bits of $m(u)$ as $m(u)|\ell$. In the case of skip lists, level $\ell$ has only one list, for each $\ell$, which contains all elements $u$ such that $m(u)|\ell=1^{\ell}$, i.e., all elements whose first $\ell$ coin flips all came up heads. Skip graphs, instead, have $2^{\ell}$ lists at level $\ell$, which we can index from 0 to $2^{\ell}-1$. Node $u$ belongs to the $j$-th list of level $\ell$ if and only if $m(u)|\ell$ corresponds to the binary representation of $j$. Hence, each node is present in one list of every level until it eventually becomes the only member of a singleton list. Without loss of generality, assume that the skip graph has sentry nodes as classical skip lists [^11] Thus, at each level $\ell$ there are $2^{\ell}$ sentinels (one for each list in such level).. Thus, the skip graph can be maintained as follows:
@@
To build the base level of the buffer skip graph we can use the same $\mathcal{O}(\log n)$ rounds approach described in Section 2.4. While, to construct the skip graph from level 0, we build an “augmented” skip list in which each node $u$ in each level $\ell$ is identified by its ID/key and the list $j$ at level $\ell$ in which $u$ appears (there are $2^{\ell}$ lists at level $\ell$). Thus, each node $u$ at some level $\ell$ can be: (i) effective in the list $j$ and fill-in for the remaining $i\neq j$, for $j,i\in[1,2^{\ell}]$; or, (ii) fill-in in all the lists $j\in[1,2^{\ell}]$. Finally, such augmented skip list is transformed into a skip graph by running (in parallel) Algorithm 2 in Section 2.3 on each level $\ell$ for which there exists at least one fill-in node.
@@
Corollary 2.28.
The maintenance cycle described in Algorithm 1 can be adapted to support skip graphs. In particular, the resulting data structure remains resilient to churn at a rate of up to $\mathcal{O}(n/\log n)$ per rounds, and all operations complete within $\mathcal{O}(\log n)$ rounds with high probability, using at most $\mathcal{O}(\text{polylog}{(n)})$ messages per node per round.
@@ 2.8 Extending our approach to deal with multiple keys on each node
Throughout the paper we assumed that each node in the network posses exactly one element of the data structure. This assumption can be relaxed to deal with multiple keys on each node. Indeed, in the case in which in the network there are $t\cdot n$ keys, where $t=\text{polylog}{(n)}$, all the techniques described above allow to maintain the data structure in the presence of the same adversarial churn rate.
@@
Given a network with $n$ nodes in which each vertex $v$ possesses $t=\text{polylog}{(n)}$ elements in the skip list. Then maintenance protocol requires $\mathcal{O}(\log n)$ rounds to build and maintain a resilient skip list that can withstand heavy adversarial churn at a churn rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round.
The corollary follows from the fact that the adversary can assign the $t$ elements to each node in the network in such a way that each of their elements will be in a different cohesive group. Indeed, the first three phases of the maintenance cycle (i.e., Delete, Creation and Merge) will require $\mathcal{O}(\log(t\cdot n))$ rounds and for $t=\text{polylog}{(n)}$ we have that the running time of the maintenance cycle is $\mathcal{O}(\log(\text{polylog}{(n)}\cdot n))=\mathcal{O}(\log\log^{k}n+\log n)=\mathcal{O}(\log n)$, where $k>0$. Furthermore, to deal with larger $t$ the protocol must be slightly adjusted to take care of the higher amount of keys per node. Here, we do not delve into this problem since it boils down to an implementation detail that can be taken care of while implementing the protocol on real world P2P Networks.
3 Implications and Generalization of the Maintenance Framework
We showed, for the first time, that complex pointer-based data structures can be maintained efficiently in the DNC model under near-linear adversarial churn. Our four-phase maintenance cycle provides the foundation for a general approach to maintaining overlays in highly dynamic environments. In Section 3.1 we propose some possible adaptations of our results to other distributed computing problems in the DNC model. In Section 3.2 we rephrase our results in terms of general framework for any distributed data structure that satisfies some specific properties. Finally, we propose a classification of distributed data structures in the DNC model that depends on the churn rate that they can tolerate.
3.1 Applications of the Maintenance Framework
We describe several case studies illustrating how our maintenance protocol could allow other distributed data structures to be maintained in highly dynamic networks.
Maintaining expander graphs.
Minimum Spanning Tree Maintenance.
Maintaining Skip Graphs.
Another implication is the construction and maintenance of other more sophisticated data structures like skip graphs [3, 39, 47]. All these data structures are not resilient to churns, and their maintenance protocols are not fast enough to recover the data structure after the failure of some nodes. Indeed, these protocols (see [3, 39, 47]) might need $\mathcal{O}(n)$ rounds to rebuild the skip-graph and they strictly require no additional churn to be able to fix the data structure. Our maintenance protocol overcomes these problems and provides a $\mathcal{O}(\log n)$ rounds skip graph constructing protocol and a technique able to repair them in $\mathcal{O}(\log n)$ in the presence of an almost linear churn of $\mathcal{O}(n/\log n)$ at every round. Moreover, our maintenance mechanism allows for the users to query the data structure while being maintained.
Beyond Distributed Computing.
We believe that some ideas from our results could be used in the centralized batch parallel setting to quickly insert batches of new elements in skip lists (or skip graphs) data structures (see e.g.,[82]) and in the fully dynamic graph algorithms settings (see for example the survey [43]) to perform fast updates of fully dynamic data structures. Indeed, in principle (provided the right amount of parallelism), our deletion and merge algorithms could be implemented in a parallel (centralized) setting and used to speed-up all kinds of computations involving these specific data structures.
3.2 A General Framework for Churn-Resilient Structures
On a broader picture, the above applications can be unified under a single general framework for maintaining pointer-based data structures under adversarial churn. Given a data structure $\mathcal{D}$, we can “abstract away” the ideas in Algorithm 1 and obtain the following abstract maintenance cycle for a generic data structure $\mathcal{D}$:
  1. 1.
    Delete Phase: Identify and remove corrupted, outdated, or disconnected regions of $\mathcal{D}$. Deletion ensures that inconsistencies caused by churn do not propagate through the structure.
  2. 2.
    Creation Phase: Organize the set of newly arrived nodes into a provisional structure $\mathcal{B}$ (the buffer data structure). The goal is to prepare these nodes for integration, typically by arranging them into a sorted or partially structured form.
  3. 3.
    Merge Phase: Integrate the buffer network $\mathcal{B}$ into the main structure $\mathcal{D}$ using a distributed merge protocol. This phase reconstructs the structure while respecting existing invariants.
  4. 4.
    Update Phase: Perform any necessary local corrections, including pointer rebalancing, level adjustments, or redundancy restoration, to finalize the integration.
Provided that we have $\mathcal{O}(T)$-round distributed algorithms for each phase of the above abstract maintenance cycle, we can maintain $\mathcal{D}$ against an adversarial churn rate of $\mathcal{O}(n/T)$ per round.
Theorem 3.1.
Let $\mathcal{D}$ be a distributed pointer-based data structure maintained using our four-phase cycle, and let $T$ be the maximum number of rounds needed for any phase. Then $\mathcal{D}$ can tolerate an adversarial churn rate of up to $\mathcal{O}(n/T)$ nodes per round, while preserving global correctness for at least $n^{c}$ rounds w.h.p., for any fixed constant $c>0$.
The proof of the theorem follows the same steps as the proof of Theorem 1.2 by considering the $\mathcal{O}(T)$-round algorithms for the deletion, buffer creation, merge, and update phases for the specific data structure $\mathcal{D}$.
Classes of Churn-Resilient Data Structures.
Our abstract maintenance cycle allows to define classes of churn-resilient data structures in the DNC model, where each class is characterized by the churn rate per round (thus the maintenance cycle runtime) that the data structure can tolerate while still supporting efficient update and query operations. Formally, let $t$ be the maintenance cycle run time function. We say that a distributed data structure $\mathcal{D}$ is $t$-maintainable if it can be successfully maintained by a $t$-round maintenance cycle. {observation} Under the above general framework, every $t$-maintainable distributed data structure tolerates an adversarial churn rate of $\mathcal{O}(n/t)$ per round. In this paper, we showed that Skip-Lists (Section 2) and Skip-Graphs (Section 2.7) are both $\mathcal{O}(\log n)$-maintainable and that are robust against and oblivious adversarial churn rate of $\mathcal{O}(n/\log n)$ per round. We conjecture that our bounds are tight, in the sense that $\log n$ is also a lower bound for the maintenance cycles for these data structures. That is because, in order to beat the $\Omega(\log n)$ barrier we would need to solve distributed sorting faster than in $\log n$ rounds while maintaining $(\alpha,\beta)$-dynamic resource competitiveness.
This classification also raises a number of intriguing open problems. For example, it remains unclear whether there exist distributed data structures in the DNC model that are $\log\log n$-maintainable, and more generally, how to characterize the precise boundaries between maintainability classes. A key challenge is to establish lower bounds on the maintenance cycle time for fundamental distributed data structures. Do entirely new data structures need to be designed to exploit faster maintenance cycles? We believe that our formulation of maintainability classes in the DNC model opens up a rich landscape for further exploration.
4 Related Works
@@ 5 Concluding remarks and discussion
In this work we proposed the first churn resilient skip list that can tolerate a heavy adversarial churn rate of $\mathcal{O}(n/\log n)$ nodes per round. The data structure can be seen as a four networks architecture in which each network plays a specific role in making the skip list resilient to churns and keeping it continuously updated. Moreover, we provided efficient $\mathcal{O}(\log n)$ rounds resource competitive algorithms to (i) delete a batch of elements from a skip list (ii) create a new skip list and, (iii) merge together two skip lists. This last result is the first algorithm that can merge two skip lists (as well as a skip list and a batch of new nodes) in $\mathcal{O}(\log n)$ rounds w.h.p.. We point out that these algorithms can be easily adapted to work on skip graphs [3, 39, 47].
In a broader sense, our technique is general and can be seen as a framework to maintain any kind of distributed data structure despite heavy churn rate. The only requirement is to devise efficient delete, buffer creation, merge, and update algorithms for the designated data structure. Furthermore, this allows us to define complexity classes for the maintenance of distributed data structure in the Dynamic Networks with Churn Model. Indeed, in this work we showed that skip list and skip graphs belong to the class of data structure that can tolerate a churn rate of $\mathcal{O}(n/\log n)$ per round.
An additional contribution of our work is the improvement on the $\mathcal{O}(\log^{3}n)$ rounds state-of-the-art technique for sorting in the $\text{NCC}_{0}$ model. We show how to sort $n$ elements (despite a high churn rate) in $\mathcal{O}(\log n)$ rounds using results from sorting network theory. However, given the impracticability of the AKS sorting network, our result is purely theoretical. In practice, it could be easily implemented using Batcher’s network instead of the AKS one. This change would slow down the bootstrap phase and the maintenance cycle to $\mathcal{O}(\log^{2}n)$ rounds. The churn rate that we can tolerate will only drop down to $O(n/\log^{2}n)$.